Opened 8 years ago

Closed 3 years ago

#384 closed flight processing (fixed)

EUFAR10/08 (ValCalHyp), flight day 301/2010, ValCalHyp

Reported by: iopa Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data location: ~arsf/arsf_data/2010/flight_data/archive_overflow/EUFAR10_08-2010_301_ValCalHyp

Data arrived from ARSF via network transfer on 4th November 2010.

Scientific objectives:

  • addressing atmospheric correction performance and related data quality

as a function of sun elevation

  • validity of change detection approaches under varying conditions

(different times, sensors, sun elevation)

Priority: a4h

PI: Eyal Ben-Dor

Sensors:

  • Eagle (18/04/2011)
  • Hawk (18/04/2011)
  • Leica LIDAR (19/09/2011)
  • RCD Digital Photography (18/04/2011)

Change History (33)

comment:1 Changed 8 years ago by iopa

Note that eagle flightline 6 had invalid GPS Start time and position. They have been replaced with time/position of hawk flightline 6.

comment:2 Changed 8 years ago by mggr

When I was at Toulouse, the PI requested a quick transfer of the raw data (E & H ; may ask for nav later). Please can we put this up on the FTP site? When you send the data arrival notification, mention we're doing this.

It'd be worth also informing him of the wavelengths in the Hawk .hdr file being off by 20nm due to the calibration issue (hopefully to be resolved in the next couple of weeks).

comment:3 Changed 8 years ago by iopa

  • Description modified (diff)

comment:4 Changed 8 years ago by benj

  • Description modified (diff)

comment:5 Changed 8 years ago by iopa

  • Description modified (diff)

comment:6 Changed 8 years ago by anch

Ioannis put the RAW data on FTP on Nov 4th when the data arrived.

Taking hyperspectral processing.

comment:7 Changed 8 years ago by anch

Creating ASTER DEM to process data as the DEM we have doesn't cover the whole flight area.

NASA WIST services are currently down (So I can't download the ASTER tiles) I'll try this again later.

comment:8 Changed 8 years ago by anch

New DEM created at ~arsf/dems/southfrance (ASTER tiles in ~arsf/ASTER) and hyperspectral has been submitted for processing.

comment:9 Changed 8 years ago by anch

Hyperspectral ready for delivery check

comment:10 Changed 8 years ago by anch

Processing cam from Raw to tif

comment:11 Changed 8 years ago by anch

Processing camera

comment:12 Changed 8 years ago by emca

Hyperspectral delivery check complete.
All data is good, ready to deliver.

comment:13 Changed 8 years ago by anch

Camera submitted for delivery check

comment:14 Changed 8 years ago by anch

  • Summary changed from EUFAR10/08, flight day 301/2010, ValCalHyp to EUFAR10/08 (ValCalHyp), flight day 301/2010, ValCalHyp

comment:15 Changed 8 years ago by emca

Delivery checked camera data. All looks good, ready to deliver.

comment:16 Changed 8 years ago by anch

  • Description modified (diff)

Camera and hyperspectral delivered to PI.

Hyperspectral delivered with HYQUAPRO extras:

  • ISO19115 XML
  • Full and flattened mask files
  • Nav quality vector

comment:17 Changed 8 years ago by knpa

SCTs

eagle hawk
1 0.00 0.00
2 0.00 0.00
3 0.02 0.00
4 0.03 0.03
5 0.03 0.00
6 0.03 0.00
7 0.04 0.00
8 0.03 0.00
9 0.04 0.00
10 0.03 0.03
11 0.00 0.00
12 0.00 0.00
13 0.00 0.00
14 0.03 0.00
15 0.03 0.00

comment:18 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Milestone set to 2010 data processing completion
  • Owner set to knpa
  • Status changed from new to assigned

Project has been r-synced.

Beginning archiving.

comment:19 Changed 8 years ago by knpa

Tarball created, awaiting NEODC download (27/05/11)

comment:20 Changed 8 years ago by knpa

Backed up by NEODC. Moved to archive disk.

comment:21 Changed 8 years ago by emca

Starting LiDAR processing

comment:22 Changed 8 years ago by emca

Processed IPAS data. Had to process in two parts to get acceptable results.

According to the log sheet there was a problem with the LiDAR part way through the data collection, and there is one line (UNMIX) that does not have LiDAR data. This may be the cause of the problems with the navigation.

comment:23 Changed 8 years ago by emca

There is no LiDAR data for line STS1 (as well as line UNMIX mentioned above).

Lines SVC1 - SVC6 have very poor returns (< 1%) - these lines are over water and are mainly just noise.

In summary:
The first 8 lines have little or no useful data.
Lines 9-15 contain good data (STS2 - STS8).

comment:24 Changed 8 years ago by emca

LiDAR data ready for delivery check at
/users/rsg/airborne/workspace/EUFAR10_08-2010_301_ValCalHyp/delivery/20110815

comment:25 Changed 7 years ago by adbe

LiDAR ready to deliver

comment:26 Changed 7 years ago by emca

Contacted PI to see if he would prefer the data on a disk or by FTP (the last disk that was sent to him had some problems clearing customs).


comment:27 Changed 7 years ago by emca

  • Description modified (diff)

LiDAR data uploaded to ftp site 19/09/2011 as requested by PI.

comment:28 Changed 7 years ago by emca

LiDAR data needs to be archived at NEODC. The rest of the data has been archived previously.

Have not r-synced the project as it is now at /users/rsg/arsf/arsf_data/archived... and don't want to mixed archived and non-archived data.

Reassigning ticket to knpa for archiving.

comment:29 Changed 7 years ago by knpa

Created tarball of LiDAR delivery and made available to NEODC (14/12/2011)

comment:30 Changed 7 years ago by knpa

NEODC have ingested the above.

comment:31 Changed 6 years ago by knpa

Have moved LiDAR delivery to archive disk.

comment:32 Changed 6 years ago by knpa

Removed workspace version (original processing).

Leaving ticket open until reprocessing is archived.

comment:33 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Reprocessed hyperspectral data now available from NEODC (http://browse.ceda.ac.uk/browse/neodc/arsf/2010/EUFAR10_08/EUFAR10_08-2010_301_ValCalHyp/hyperspectral-reprocessed) - closing ticket.

Note: See TracTickets for help on using tickets.